# Generalizing matrix representations to fully heterochronous ranked tree shapes

### v3

@@ 3 Theorems
Items 2 and 3 are immediate by definitions. While item 1 may also appear obvious, we do not know a priori that $L_{f}(N,M+1)\leq U_{f}(N,M+1)$. In fact, this is the major claim in justifying that such sequences extend. However, when doing so we may freely use bounds for $L_{f}(i,j)$ and $U_{f}(i,j)$ at previous indices with $0\leq j\leq i-2$.
Suppose M<N2M<N-2. Our first goal is to verify that
$f_{N,M}\leq f_{N-1,M}\leq f_{N-1,M+1}.$ (4)
By definition 1.2, $f_{N,M}\leq U_{f}(N,M)\leq f_{N-1,M}$. We handle the remaining bound in three cases. First note that by definitions 1.3 and 1.1,
$f_{N-2,N-3}=f_{N-3,N-3}-1=(f_{N-2,N-2}\pm 1)-1.$
When M=N3M=N-3 and $f_{N-2,N-3}=f_{N-2,N-2}$, by definition 1.3,
$f_{N-1,M}=f_{N-1,N-3}=f_{N-2,N-2}-1=f_{N-1,N-2}=f_{N-1,M+1}.$
When M=N3M=N-3 and $f_{N-2,N-3}=f_{N-2,N-2}-2$, definitions 1.2 and 1.3 give
$<f_{N-1,M+1}.$
Lastly, when M<N3M<N-3, by definition 1.2,
$f_{N-1,M+1}\geq L_{f}(N-1,M+1)\geq f_{N-1,M}.$
Therefore (4) is true.
We consider the two possible values for $U_{f}(N,M+1)$. If $U_{f}(N,M+1)=f_{N-1,M+1}$, then $f_{N,M}\geq f_{N-1,M}$ and so in fact $f_{N,M}=f_{N-1,M}$. Therefore,
@@
Furthermore, if we additionally have $f_{N-1,M+1}=f_{N-1,N-1}$, then $f_{N,M}<f_{N-1,M+1}$, as $f_{N,M}=f_{N-1,M+1}$ implies $f_{N-1,M}=f_{N-1,N-1}$, which yields the contradiction $f_{N,M}=f_{N-1,N-1}-1$. Specifically, with the additional assumption that $f_{N-1,M+1}=f_{N-1,N-1}$, we have $L_{f}(N,M+1)=f_{N-1,N-1}-1$.
When instead $U_{f}(N,M+1)=f_{N,M}+f_{N-1,M+1}-f_{N-1,M}<f_{N-1,M+1}$, we have $f_{N,M}<f_{N-1,M}\leq f_{N-1,M+1}$, so that
@@
$\leq f_{N-1,M+1}+f_{N,M}-f_{N-1,M}=U_{f}(N,M+1)$
Furthermore, if we additionally have $f_{N-1,M+1}=f_{N-1,N-1}$, then $L_{f}(N,M+1)=f_{N-1,N-1}-1$. This establishes item 1.
Lastly, we verify item 4. We have
@@
If $f_{N,0}=f_{N,N}$, then $f_{N,0}\geq 1$ and so $L_{f}(N+1,0)=f_{N-1,N-1}-1$. ∎
@@
Suppose $F$ is an F-matrix. The sequence $F_{i,j}$ immediately satisfies all conditions of an (n,2n3,2n3)(n,2n-3,2n-3) F-sequence, except possibly the condition when $F_{i-1,j}=F_{i-1,i-1}$. If $F_{i-1,j}=F_{i-1,i-1}$ and $0\leq j\leq i-1$, then by conditions 2, 1, and 33(b) of Theorem 2,
@@
Let us emphasize that Proposition 1 provides the rules used to construct the entries, one at a time, of an F-matrix. When following these rules, there is no chance of entering an invalid state that requires backtracking to previously selected entries. Furthermore, when constructing an entry, we need only consider values at four previous entries (the entries directly to the left, directly above, and directly to the above-left, as well as the previous diagonal entry). As such we have an efficient process to determine all F-matrices of a given size and so all ranked tree shapes on a given number of leaves. Note it is a straightforward process to turn an F-matrix into an E-matrix, and to turn an E-matrix into a ranked tree shape.
@@ 4 Sampling Schemes
Another consequence of the bijection between ranked tree shapes and full-cherry trees is that we can recursively count the number of fully heterochronous ranked tree shapes via a standard recursion involving root-splitting (Steel-2016) to obtain the following proposition.
@@
Let $K_{n}$ denote the number of fully heterochronous ranked tree shapes with $n$ leaves. Then $K_{n}$ satisfies the following initial conditions and recursion,
$K_{1}=K_{2}=1,\qquad K_{n}=\frac{1}{2}\sum_{\ell=1}^{n-1}\binom{2n-2}{2\ell-1}K_{\ell}K_{n-\ell},$ (5)
@@
As the initial conditions are trivial, we assume $n\geq 3$. To prove the recursion, we use that $K_{n}$ is also the number of full-cherry trees with $2n$ leaves. The full-cherry trees with $2n$ leaves may be constructed as follows. Select two full-cherry trees $T_{1}$ and $T_{2}$, where $T_{1}$ has $2\ell$ leaves and $T_{2}$ has $2(n-\ell)$ leaves with $0<\ell<n$; extend the total orderings of the internal nodes of $T_{1}$ and $T_{2}$ to a common total ordering; join $T_{1}$ and $T_{2}$ with a new root node whose children are the roots of $T_{1}$ and $T_{2}$. Since $T_{1}$ has $2\ell-1$ internal nodes and $T_{2}$ has $2(n-\ell)-1$ internal nodes, there are exactly $\binom{2n-2}{2\ell-1}$ ways to extend to a common ordering. The sum in (5) corresponds to this construction, where the factor $\frac{1}{2}$ accounts for double counting due to constructing full-cherry trees from ordered pairs $(T_{1},T_{2})$ rather than sets $\{T_{1},T_{2}\}$. ∎
We note that this recursion agrees with the recursion for strictly ordered binary trees by Poupard poupard1989. Indeed, Poupard’s strictly ordered binary tree is a different name for the fully heterochronous ranked tree shape. Using this recursion and an argument by generating functions, Poupard showed that the number of strictly ordered binary trees with $n$ leaves is equal to the $n^{th}$ reduced tangent number (see (1)).
In this section we introduce three methods for sampling fully heterochronous ranked tree shapes. The first method is a coalescent model inspired by the bijection with full-cherry trees (Palacios2015, Proposition 2). This model is “bottom-up” in the sense that the generating process starts with one cherry node and adds cherries and merges cherries one by one until the root. The second method, in contrast, is “top-down”: it starts with the root and sequentially selects edges to bifurcate or to sample (terminate) as time moves forward. This method utilizes the Catalan diagonal structure of the $\mathbf{F}$-matrix. The last model generates one entry at a time sequentially along the $\mathbf{F}$-matrix via Bernoulli probabilities. This last model can be specialized to a class of Beta-splitting models.
@@ 4.1 Coalescent model
Figure 3: An example of the coalescent jump chain with states $A_{i}$ and corresponding transition probabilities.
In order to compute the probability of a fully heterochronous ranked tree shape in terms of its $\mathbf{F}$-matrix $F$, we need to determine the number of unmerged leaves and unmerged ranked nodes of the fully-cherry tree from $F$. For the number of unmerged ranked nodes, we have
@@ 4.2 Diagonal “top-down” model
It is well-known that the number of Dyck sequences of length 2n22n-2 is the Catalan number $C_{n-1}=\frac{1}{n}\binom{2(n-1)}{n-1}$ Stanley-2012. Therefore, the number of possible diagonals in the F-matrices of fully heterochronous trees with $n$ leaves is the Catalan number $C_{n-1}$. ∎
An algorithm to sample Dyck paths from $(1,0)$ to (n1,n1)(n-1,n-1) that has $O(n)$ complexity was proposed by Devroye1999. The algorithm proceeds sequentially starting from (1,0); at any point $(i,j)$ in the partially formed Dyck path, we move to the right with probability $N(i+1,j)/N(i,j)$, where
@@
Notice that $P(F)$ depends solely on the fixed diagonal $\{F_{k,k}\}_{k=0}^{2n-3}$, so all fully heterochronous ranked tree shapes with the same diagonal have the same probability. This concludes the proof. ∎
@@ 4.3 A Bernoulli splitting model
We further note that in determining valid values, we need at most four previous values rather than all previous values. Let $\mathcal{F}_{i,j}\mid F_{<(i,j)}$ denote the set of possible values that $F_{i,j}$ can take conditionally of previous values, then given real numbers $p_{i,j}\in(0,1)$, with $(i,j)$ ranging over the non-trivial entries, entry $F_{i,j}$ is $L_{F}(i,j)=\min(\mathcal{F}_{i,j}\mid F_{<(i,j)})$ with probability $p_{i,j}$ and $U_{F}(i,j)=\max(\mathcal{F}_{i,j}\mid F_{<(i,j)})$ with probability $1-p_{i,j}$. We set
$P(F_{i,j}\mid F_{<(i,j)},p_{i,j})\\ =\delta_{L_{F}(i,j)=U_{F}(i,j)}+(1-\delta_{L_{F}(i,j)=L_{F}(i,j)})\,p_{i,j}^{\delta_{F_{i,j}=L_{F}(i,j)}}(1-p_{i,j})^{\delta_{F_{i,j}=U_{F}(i,j)}}$
The joint probability of an F-matrix conveniently telescopes, as
@@
There are four F-matrices for fully heterochronous ranked tree shapes with three leaves:
$F^{0}$ $=\begin{pmatrix}2&0&0&0\\ 1&1&0&0\\ 0&0&2&0\\ 0&0&1&1\end{pmatrix}$, $F^{1}$ $=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 0&2&2&0\\ 0&1&1&1\end{pmatrix}$,
$F^{2}$ $=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 1&2&2&0\\ 0&1&1&1\end{pmatrix}$, $F^{3}$ $=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 1&2&2&0\\ 1&1&1&1\end{pmatrix}$.
@@
Inspired by the Beta-splitting model aldous1996probability; sainudiin2016beta, we can sample the Bernoulli probabilities from a Beta density $f(p_{i,j};\alpha,\beta)$, with parameters $\alpha$ and $\beta\in(0,\infty)$. The entry-wise probability is
@@
This Beta-Bernoulli model generates balanced trees when $\alpha\gg\beta$ and unbalanced otherwise. The mean of Beta($\alpha,\beta$) is $\frac{\alpha}{\alpha+\beta}$, so when $\alpha\gg\beta$, the Bernoulli probabilities of choosing the smaller admissible value $L_{F}$ tends to be higher. This results in lineages surviving less, which gives more balanced trees. This can be appreciated in Figure 6 obtained from simulations.

### v4

@@ 3 Theorems
The proof is to verify that the extended sequences satisfy Definition 2. We do so in the order of the four cases of the size of $M$ relative to $N$. However, we do not split each case into the subcases of the statement of the proposition.
We begin with Item 1, where we assume M<N2M<N-2. This is the lengthiest case to prove. Since the sequence is not extended with a diagonal entry, the extended sequence immediately satisfies Definition 2.1. To verify the extended sequence satisfies Definitions 2.2 and 2.3, we show that $L_{f}(N,M+1)\leq U_{f}(N,M+1)$ and that if $f_{N-1,M+1}=f_{N-1,N-1}$ then $L_{f}(N,M+1)=f_{N-1,N-1}-1$.
At the outset, we do not know that $L_{f}(N,M+1)\leq U_{f}(N,M+1)$; we only have that $L_{f}(i,j)\leq f_{i,j}\leq U_{f}(i,j)$ for $0\leq j\leq i-2$ with $(i,j)$ before (N,M+1){(N,M+1)} in lexicographical order. Recall that
$U_{f}(N,M+1)=\min(f_{N-1,M+1},f_{N,M}+f_{N-1,M+1}-f_{N-1,M}).$
We consider the two possible values for $U_{f}(N,M+1)$.
Suppose $U_{f}(N,M+1)=f_{N-1,M+1}$, which means $f_{N,M}\geq f_{N-1,M}$. However, $f_{N,M}\leq f_{N-1,M}\leq f_{N-1,M+1}$ by Lemma 1, so in fact $f_{N,M}=f_{N-1,M}$, and therefore,
@@
What remains is to show that if $f_{N-1,M+1}=f_{N-1,N-1}$ then $L_{f}(N,M+1)=f_{N-1,N-1}-1$. Suppose $f_{N-1,M+1}=f_{N-1,N-1}$. We claim that $f_{N,M}<f_{N-1,M+1}$. For if not, meaning $f_{N,M}=f_{N-1,M+1}=f_{N-1,N-1}$, then $f_{N-1,M}=f_{N-1,M+1}=f_{N-1,N-1}$ by Lemma 1 and so $f_{N,M}=f_{N-1,N-1}-1$ by Definition 2.3, which is a contradiction. So if $f_{N-1,M+1}=f_{N-1,N-1}$ then $L_{f}(N,M+1)=f_{N-1,N-1}-1$. In summary, under the assumption that $U_{f}(N,M+1)=f_{N-1,M+1}$, the extended sequence satisfies Definition 2.
Suppose instead $U_{f}(N,M+1)=f_{N,M}+f_{N-1,M+1}-f_{N-1,M}<f_{N-1,M+1}$, which means $f_{N,M}<f_{N-1,M}$. However, $f_{N-1,M}-1\leq f_{N,M}$ by Definition 2.2, so in fact $f_{N,M}=f_{N-1,M}-1$, whereas $f_{N-1,M}\leq f_{N-1,M+1}$ by Lemma 1. Therefore
@@
$=f_{N-1,M+1}+f_{N,M}-f_{N-1,M}=U_{f}(N,M+1).$
If $f_{N-1,M+1}=f_{N-1,N-1}$, then trivially $L_{f}(N,M+1)=f_{N-1,N-1}-1$. In summary, under the assumption that $U_{f}(N,M+1)\not=f_{N-1,M+1}$, the extended sequence satisfies Definition 2. This completes the proof of Item 1.
For Item 2, we note that $f_{N-1,N-1}-1$ is exactly the value of $f_{N,N-1}$ specified by Definition 2.3 . Since the sequence is extended by a subdiagonal entry, the extended sequence immediately satisfies Definitions 2.1 and 2.2.
For Item 3, we note that these values of $f_{N,N}$ are exactly those specified by Definition 2.1. Since the sequence is extended by a diagonal entry, the extended sequence immediately satisfies Definitions 2.2 and 2.3.
Last, we verify Item 4. Since the sequence is not extended with a diagonal entry, the extended sequence immediately satisfies Definition 2.1. As with Item 1, we show that the extended sequence satisfies Definitions 2.2 and 2.3 by showing that $L_{f}(N+1,0)\leq U_{f}(N+1,0)$ and that if $f_{N,0}=f_{N,N}$ then $L_{f}(N+1,0)=f_{N,N}-1$. We have
@@
If $f_{N,0}=f_{N,N}$, then $f_{N,0}\geq 1$ and so $L_{f}(N+1,0)=f_{N,0}-1=f_{N-1,N-1}-1$.
Example 5.
Using Proposition 1 above, we can avoid the conflict in Example 3. To extend the $(3,3,0)$ F-sequence
$\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 0&2&2&0\\ 0&*&*&*\end{pmatrix}$,
Item 1(a) tells us $f_{3,1}$ must be 11, and the resulting $(3,3,1)$ F-sequence is
$\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 0&2&2&0\\ 0&1&*&*\end{pmatrix}$.
Notably, the illegal value $f_{3,2}=2$ is no longer allowed.
@@
Suppose $F$ is an F-matrix. The sequence $F_{i,j}$ immediately satisfies all conditions of an (n,2n3,2n3)(n,2n-3,2n-3) F-sequence, except possibly the condition when $F_{i-1,j}=F_{i-1,i-1}$ in Definition 2.3. If $F_{i-1,j}=F_{i-1,i-1}$ and $0\leq j\leq i-1$, then by Conditions 2, 1, and 33(b) of Theorem 2,
@@
Let us emphasize that Proposition 1 provides the rules used to construct the entries, one at a time, of an F-matrix. When following these rules, there is no chance of entering an invalid state that requires backtracking to previously selected entries. Furthermore, when constructing an entry, we need only consider values at four previous entries (the entries directly to the left, directly above, and directly to the above-left, as well as the previous diagonal entry). As such we have an efficient process to determine all F-matrices of a given size and so all ranked tree shapes on a given number of leaves. Note it is a straightforward process to turn an F-matrix into an E-matrix using the equations in (2), and to turn an E-matrix into a ranked tree shape. This method of enumeration is illustrated for the $4\times 4$ F-matrices in Figure 2, with the corresponding fully heterochronous ranked tree shapes drawn in Figure 3.
{forest}
Figure 2: The iterative construction of all $4\times 4$ F-matrices. The positions that admit two possible values are indicated by the * sign, with the two downward edges representing each of the values. Intermediate steps for positions with a single possible value are omitted. The leaves ($F^{0},F^{1},F^{2},F^{3}$) enumerate all $4\times 4$ F-matrices.
Figure 3: The 33-leaf fully heterochronous ranked tree shapes corresponding to each of the $4\times 4$ F-matrices.
@@ 4 Sampling Schemes
Another consequence of the bijection between ranked tree shapes and full-cherry trees is that we can recursively count the number of fully heterochronous ranked tree shapes via a standard recursion involving root-splitting (Steel, 2016, Chapter 2.1) to obtain the following proposition.
@@
The number of fully heterochronous ranked tree shapes with $n$ leaves, $\left|\mathcal{T}^{*}_{n}\right|$, satisfies the following initial conditions and recursion,
$\left|\mathcal{T}^{*}_{1}\right|=\left|\mathcal{T}^{*}_{2}\right|=1,\qquad\left|\mathcal{T}^{*}_{n}\right|=\frac{1}{2}\sum_{\ell=1}^{n-1}\binom{2n-2}{2\ell-1}\left|\mathcal{T}^{*}_{\ell}\right|\left|\mathcal{T}^{*}_{n-\ell}\right|,$ (4)
@@
As the initial conditions are trivial, we assume $n\geq 3$. To prove the recursion, we use that $\left|\mathcal{T}^{*}_{n}\right|$ is also the number of full-cherry trees with $2n$ leaves. The full-cherry trees with $2n$ leaves may be constructed as follows. Select two full-cherry trees $T_{1}$ and $T_{2}$, where $T_{1}$ has $2\ell$ leaves and $T_{2}$ has $2(n-\ell)$ leaves with $0<\ell<n$; extend the total orderings of the internal nodes of $T_{1}$ and $T_{2}$ to a common total ordering; join $T_{1}$ and $T_{2}$ with a new root node whose children are the roots of $T_{1}$ and $T_{2}$. Since $T_{1}$ has $2\ell-1$ internal nodes and $T_{2}$ has $2(n-\ell)-1$ internal nodes, there are exactly $\binom{2n-2}{2\ell-1}$ ways to extend to a common ordering. The sum in (4) corresponds to this construction, where the factor $\frac{1}{2}$ accounts for double counting due to constructing full-cherry trees from ordered pairs $(T_{1},T_{2})$ rather than unordered pairs $\{T_{1},T_{2}\}$. ∎
We note that this recursion agrees with the recursion for strictly ordered binary trees by Poupard Poupard (1989), although her proof does not involve full-cherry trees. Indeed, Poupard’s strictly ordered binary tree is a different name for the fully heterochronous ranked tree shape. We restate the result to emphasize the value of working with a full-cherry tree, which will be the building block of the coalescent model introduced below. Using this recursion and an argument by generating functions, Poupard showed that the number of strictly ordered binary trees with $n$ leaves is equal to the $n^{th}$ reduced tangent number (see (1)). Table 2 compares the cardinality of the space of isochronous ranked tree shapes and that of fully heterochronous ranked tree shapes for varying numbers of leaves.
Number of leaves (n)
11 22 33 44 55 66 77 88 99 $10$
$\left|\mathcal{T}_{n}\right|$ 11 11 11 22 55 $16$ $61$ $272$ $1385$ $7936$
$\left|\mathcal{T}^{*}_{n}\right|$ 11 11 44 $34$ $496$ $11056$ $349504$ $14873104$ $819786496$ $56814228736$
Table 2: The number of isochronous and heterochronous ranked tree shapes with $n$ leaves, denoted respectively by $\left|\mathcal{T}_{n}\right|$ and $\left|\mathcal{T}^{*}_{n}\right|$.
In this section we introduce three methods for sampling fully heterochronous ranked tree shapes. The first method is a coalescent model inspired by the bijection with full-cherry trees (Palacios et al., 2015, Proposition 2). This model is “bottom-up” in the sense that the generating process starts with one cherry node and adds cherries and merges cherries one by one until the root. The second method, in contrast, is “top-down”: it starts with the root and sequentially selects edges to bifurcate or to sample (terminate) as time moves forward. This method utilizes the Catalan diagonal structure of the $\mathbf{F}$-matrix. The last model generates one entry at a time sequentially along the $\mathbf{F}$-matrix via Bernoulli probabilities. This last model can be specialized to a class of Beta-splitting models.
@@ 4.1 Coalescent model
Figure 5: An example of the coalescent jump chain with states $A_{i}$ and corresponding transition probabilities. The probability of the ranked tree shape is $\prod_{k=1}^{3}P(A_{k}|A_{k+1})=1\times\frac{1}{2}\times 1=\frac{1}{2}$.
In order to compute the probability of a fully heterochronous ranked tree shape in terms of its $\mathbf{F}$-matrix $F$, we need to determine the number of unmerged leaves and unmerged ranked nodes of the full-cherry tree from $F$. For the number of unmerged ranked nodes, we have
@@ 4.2 Diagonal “top-down” model
It is well-known that the number of Dyck sequences of length 2n22n-2 is the Catalan number $C_{n-1}=\frac{1}{n}\binom{2(n-1)}{n-1}$ Stanley (2012). Therefore, the number of possible diagonals in the F-matrices of fully heterochronous trees with $n$ leaves is the Catalan number $C_{n-1}$. ∎
An algorithm to sample Dyck paths from $(1,0)$ to (n1,n1)(n-1,n-1) that has $O(n)$ complexity was proposed by Devroye et al. (1999). The algorithm proceeds sequentially starting from (1,0); at any point $(i,j)$ in the partially formed Dyck path, we move to the right with probability $N(i+1,j)/N(i,j)$, where
@@
Notice that $P(F\mid\{F_{k,k}\}_{k=0}^{2n-3})$ depends solely on the fixed diagonal $\{F_{k,k}\}_{k=0}^{2n-3}$, so all fully heterochronous ranked tree shapes with the same diagonal have the same probability. This concludes the proof. ∎
@@ 4.3 A Bernoulli splitting model
We further note that in determining valid values, we need at most four previous values rather than all previous values. For a given non-trivial entry $F_{i,j}$, let $\mathcal{F}_{i,j}\mid F_{<(i,j)}$ denote the set of possible values that $F_{i,j}$ can take conditionally of previous values, then given real numbers $p_{i,j}\in(0,1)$, entry $F_{i,j}$ is $L_{F}(i,j)=\min(\mathcal{F}_{i,j}\mid F_{<(i,j)})$ with probability $p_{i,j}$ and $U_{F}(i,j)=\max(\mathcal{F}_{i,j}\mid F_{<(i,j)})$ with probability $1-p_{i,j}$. We set
$P(F_{i,j}\mid F_{<(i,j)},p_{i,j})\\ =\delta_{L_{F}(i,j)=U_{F}(i,j)}+(1-\delta_{L_{F}(i,j)=U_{F}(i,j)})\,p_{i,j}^{\delta_{F_{i,j}=L_{F}(i,j)}}(1-p_{i,j})^{\delta_{F_{i,j}=U_{F}(i,j)}},$
where $\delta$ is an indicator function.
Let p denote the set of Bernoulli probabilities $p_{i,j}$ with $(i,j)$ ranging over the non-trivial entries. The joint probability of an F-matrix conveniently telescopes as
@@
Figure 2 shows the four F-matrices for fully heterochronous ranked tree shapes with three leaves, and marks (with * sign) the three non-trivial positions. Recall the matrices are
$F^{0}=\begin{pmatrix}2&0&0&0\\ 1&1&0&0\\ 0&0&2&0\\ 0&0&1&1\end{pmatrix},F^{1}=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 0&2&2&0\\ 0&1&1&1\end{pmatrix},F^{2}=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 1&2&2&0\\ 0&1&1&1\end{pmatrix},F^{3}=\begin{pmatrix}2&0&0&0\\ 1&3&0&0\\ 1&2&2&0\\ 1&1&1&1\end{pmatrix}$.
@@
Inspired by the Beta-splitting model Aldous (1996); Sainudiin and Véber (2016), we can sample the Bernoulli probabilities from a Beta density $f(p_{i,j};\alpha,\beta)$, with parameters $\alpha$ and $\beta\in(0,\infty)$. The entry-wise probability is
@@
This Beta-Bernoulli model generates unbalanced trees when $\alpha\gg\beta$ and balanced otherwise. The mean of Beta($\alpha,\beta$) is $\frac{\alpha}{\alpha+\beta}$, so when $\alpha\gg\beta$, the Bernoulli probabilities of choosing the smaller admissible value $L_{F}$ (the probabilities of sampling as opposed to branching) tends to be higher. Consequently, there are typically few extant lineages at any given time, which favors caterpillar-like shapes and yields unbalanced trees. This is seen in simulations, with the total tree length indicating how long and how many branches persist before sampling (Figure 8)
Example 7.
To illustrate the differences between the three models introduced in this Section, we compute the probability of the ranked tree shape $T$ in Figure 5 under each the three models. The tree has F-matrix,
$F=\begin{pmatrix}2&0&0&0\\ 1&1&0&0\\ 0&0&2&0\\ 0&0&1&1\end{pmatrix}$.
By (5), the probability under the Coalescent model is $P_{\text{coalescent}}(T)=\frac{1}{2}$, which agrees with the direct calculations in Figure 5. Its probability under the Diagonal “top-down” model, according to Proposition 5, is
$P_{\text{diagonal}}(T)=\frac{2^{2}}{C_{2}\times 2\times 1\times 2\times 1}=\frac{1}{2}.$
Lastly, for fixed Beta parameters $\alpha$ and $\beta$, the marginal tree probability is
$P_{\text{Beta-Bernoulli}}(T)=\prod_{\text{nontrivial }(i,j)}P(F_{i,j}|F_{<(i,j)})=P(F_{1,1}|F_{<(1,1)})=\frac{\alpha}{\alpha+\beta}.$